OpenStudy (natsudragneel):

Explain how the zero property in the problem,2/3(0)=02/3=0 , is the same with any rational number.

HanAkoSolo (jamierox4ev3r):

I assume that you mean the Multiplication Property of Zero, right?

OpenStudy (natsudragneel):

Yes

HanAkoSolo (jamierox4ev3r):

So I'm having a bit of a hard time visualizing this...is this what you're trying to say? \(\Large\frac{2}{3} *~0 = 0~*\frac{2}{3}=0\)

OpenStudy (natsudragneel):

Yep exactly :)

HanAkoSolo (jamierox4ev3r):

oh okay. Since the multiplication property tells us that the product of ANY number is zero. So since \(\frac{2}{3}\) is indeed a number, it can effectively be multiplied by zero to become zero. The same goes for all other real numbers.

OpenStudy (natsudragneel):

Okay thanks so much!

HanAkoSolo (jamierox4ev3r):

no problem!